Use the information below to answer questions 17-20 A medical researcher has conducted tests on a new method of detecting hepatic (liver) tumors in heavy drinkers and wants to report his findings in the Journal of the American Medical Association. The following table shows his results. Tumor No tumor Total Positive 105 80 185 Negative 15 Total120 880 800 815 1000 17. What is the prevalence of hepatic tumors? a, 58% b. 88% c. 12% d. 85% 18. What is the sensitivity of this test? a. 58% b. 8890 c. 12% d. 85% 19. What it the false negative rate for this test? 5890 a. b, 9% 88% c. d. 9190 20. what isthe positive predictive value for this test? 2% a. b. 4390 C. 88% d, 58% 21. A Z score of 0 means that the value a. Is to the right of the mean. b. Is far from the mean c. Is close to the mean d. Is equal to the man 22. Given the alternative hypothesis (H): Consuming green tea increases a persons concentration. What would be the null hypothesis (Ho)? a. Consuming green tea decreases a persons concentration b. Consuming green tea does not increase a persons concentration. c. Consuming green tea affects a persons concentration. d. Theres not enough information. 23. Using the Standard Normal (Z) Distribution Table, find P (Z> 0.52) a. 0.6915 b. 0.3085 c. 0.3446 d. 0.6554 24. If l increase the risk of type I error (?) what would happen to the power of a study? a. Power would increase b. Power would decrease c. Power would not change d. Not enough information
Answer #1

17. The prevalance of hepatic tumors is - 120/ 1000 or 12%

18. Sensitivity = True positive/Actual positive = 105/120 = 87.5% ~ 88%

19. False negative rate = False negative/ FN + TP = 15/120 = 12.5% (Options are wrong)

20. Positive predicted value = True Positive/ Predicted Positive = 105/ 185 = 57%

21. A z value of 0 means that the score is equal to the mean